Find the coordinates of the centroid C, of the shaded area with respect to a frame of reference specified by yourself (PART2)2 3 in- 45 in УA A A у Sección Sumatorias 3. Semicirenlo 4.5 in 6 in 6 in yA A A ў Sección
Answer #1

oigin Quaskr #centroid fcircle n P 3A A. ads of cirele , 172? OAB Seetion (L) (0,0) OPR - Sechion (2 B10 OAB = 1-274in = 1-27Semiurde Thargle Ihave laken frame ef retreme Such that x-coOrdinak of shaded gion u ZERO X Semicisde N X Gin (0,0) 6in (6) 2